OpenStudy (anonymous):

Simplify. Express the quotient as a rational exponent. x^4/5 ------ x^1/3

OpenStudy (anonymous):

\[\large \frac{x^a}{x^b} = x^{a - b}\] Here you have to just subtract the powers..

OpenStudy (anonymous):

\[\frac{4}{5} - \frac{1}{3} = ??\]